OpenStudy (mollynicole2015):

Solve for v. v + 2 -------- = 2 3

OpenStudy (youngstudier):

Okay, so what do you think you should first?

OpenStudy (youngstudier):

you should do first.* Goodness. :P

OpenStudy (mollynicole2015):

Multiply v + 2 by 3?

OpenStudy (youngstudier):

Um, no..... It's v + 2 ________ = 2 3 You'd want to multiply each side by 3. So: v + 2 3 * ____________ = 2 * 3 3 This will get rid of the 3 on the left side. So what is 2 * 3?

OpenStudy (mollynicole2015):

6

OpenStudy (youngstudier):

Right. :) So now we have: v + 2 = 6 What do you think we should do next?

OpenStudy (mollynicole2015):

subtract 2 from both sides to get the variable alone.

OpenStudy (youngstudier):

Yep. :D So it would now be: v + 2 - 2 = 6 - 2 v = ?

OpenStudy (mollynicole2015):

4

OpenStudy (youngstudier):

v = 4 Good job. :D

OpenStudy (mollynicole2015):

Thank you so much! This helped a lot!

OpenStudy (youngstudier):

You're welcome! \(^_^)/

OpenStudy (mollynicole2015):

<(^-^)> I love math.

OpenStudy (youngstudier):

LOL, I do too on rare occasions XDDD ^_^

OpenStudy (mollynicole2015):

Lol.